Sphere A and B have the same initial positive charge q0. The magnitude of the repulsive electric force between the two spheres is 14.1 N when separated by a distance of 0.03 m. (a) find the initial charge q0. (b) The spheres are then allowed to touch each other and the separated. Find the electric force between them if the radius of sphere A is there times that of sphere B.

1
Expert's answer
2022-02-25T08:00:03-0500

)



14.4=9\cdot10^9\frac{q_0^2}{0.03^2}\\q_0=1.2\cdot10^{-6}\ C14.4=9⋅1090.032q02​​q0​=1.2⋅10−6 C


2)



\frac{q}{r}=\frac{q_0-q}{3r}\\q=\frac{1}{4}q_0rq​=3rq0​−q​q=41​q0​


F=9\cdot10^9\frac{3}{16}\frac{(1.2\cdot10^{-6})^2}{0.03^2}=2.7\ NF=9⋅109163​0.032(1.2⋅10−6)2​=2.7 N